# Analytic approach to boundary integrability with application to mixed-flux $AdS_{3}\times S^{3}$

### v1

@@ Abstract
Boundary integrability provides rare analytic control over field theories in the presence of an interface, from quantum impurity problems to open string dynamics. We develop an analytic framework for integrable boundaries in two-dimensional sigma-models that determines admissible reflection maps directly from the meromorphic Lax connection. Applying it to open strings on $AdS_{3}\times S^{3}$ with mixed NSNS and RR flux, we find two branches of integrable boundary conditions. One branch admits D-branes wrapping twisted conjugacy classes on $SU(1,1)\times SU(2)$, with the mixed-flux deformation encoded entirely into dynamical boundary data. At the exactly solvable WZW point these coincide with the conformal D-branes, providing a natural link to conformal perturbation theory.
@@
Field theories with boundaries play an important role in both high-energy and condensed-matter physics. In open string theory, worldsheet boundary conditions describe D-branes, while in condensed-matter and lattice systems they model defects and boundary critical phenomena. When the bulk theory is integrable, the situation becomes especially interesting, as an infinite set of conserved charges provides rare analytic control of the interacting quantum dynamics. The presence of a boundary probes this structure in a nontrivial way. Geometric symmetries are typically broken, and it is far from obvious that an infinite tower of conserved charges can survive. When they do, notions of boundary integrability have opened access to exact reflection data (see e.g. [1]) and methods for defect observables in AdS/CFT (see e.g. [2]). Identifying which physical interfaces are compatible with integrable dynamics is therefore a basic structural question. In this Letter, we address this for two-dimensional sigma-models from a novel perspective based on very simple analytical arguments.
The starting point is the one-parameter family of flat Lax connections ${\cal L}(\mathsf{x})$ representing the equations of motion of a periodic sigma-model on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$, i.e. [^1] For brevity, and where possible, we will suppress explicit time ($\tau\in\mathbb{R}_{\tau}:=\mathbb{R}$) or space ($\sigma\in S^{1}_{\sigma}:=S^{1}$ with radius 1) dependence. $\mathrm{d}{\cal L}(\mathsf{x})+{\cal L}(\mathsf{x})\wedge{\cal L}(\mathsf{x})=0$ and ${\cal L}(\sigma+2\pi;\mathsf{x})={\cal L}(\sigma;\mathsf{x})$ for every value of the spectral parameter $\mathsf{x}$. This ensures that the monodromy matrix $T_{\cal L}(2\pi,0;\mathsf{x})$, with
@@
On an interval $\sigma\in[0,\pi]$, boundary integrability requires a modification of this generator. The prototype is obtained by following Sklyanin’s construction [4, 5], where one forms a double-row monodromy matrix $T_{\text{\tiny b}}(\mathsf{x})$ by gluing $T_{\cal L}(\pi,0;\mathsf{x})$ over the interval $\sigma\in[0,\pi]$ to a reflected copy. On the lattice, Sklyanin’s approach applies to integrable models with crossing-symmetric $R$-matrices of difference form [5]. The reflection compatible with integrability then has two components: a spatial flip reversing the order of sites, and a spectral reflection $\mathsf{x}\mapsto\rho(\mathsf{x})=-\mathsf{x}$ implementing crossing symmetry. Additional boundary degrees of freedom are encoded in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ at the endpoints $\sigma_{\star}=0,\pi$, which must satisfy the Boundary Yang-Baxter Equation to preserve integrability. Translating to the sigma-model, the analogue is to apply a worldsheet parity flip $\sigma\rightarrow 2\pi-\sigma$ combined with a spectral reflection $\rho(\mathsf{x})$ that leaves the Lax connection invariant (see e.g. [4, 6, 7, 8]). One can then effectively set $T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T_{\cal L}^{-1}(\pi,0;\rho(\mathsf{x}))K_{\pi}^{-1}(\mathsf{x})T_{\cal L}(\pi,0;\mathsf{x})$ (up to automorphisms implementing other conventions) and solve the conditions for $K_{\sigma_{\star}}$ under which infinite charges are generated by $T_{\text{\tiny b}}(\mathsf{x})$.
@@
However, there are limitations to taking this prescription directly as a definition of boundary integrability for sigma-models. On the lattice, $\rho(\mathsf{x})$ is dictated by the crossing structure of a chosen $R$-matrix, and the familiar $\rho(\mathsf{x})=-\mathsf{x}$ is tied to those of difference form. For an integrable sigma-model, such quantum input is often unavailable (or only indirect through the sigma-model $S$-matrix), and the defining object is just a Lax connection ${\cal L}(\mathsf{x})$ that generically is meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$. Crucially, there need not exist a $\rho(\mathsf{x})$ that leaves ${\cal L}(\mathsf{x})$ invariant after the parity flip, and in this case there is generally not an immediate principle that selects $\rho(\mathsf{x})$ without additional structure such as e.g. quantum crossing.
@@
In this Letter, we instead determine $\rho(\mathsf{x})$ intrinsically from the analytic structure of the Lax connection, bypassing input on bulk parity or crossing invariance. [^2] Alternative Lax-based approaches via parity exist [9, *Corrigan:1996nt], but they are based on auxiliary constructions and order-by-order expansions in $\mathsf{x}$. Here we have a purely analytic approach, which moreover admits a natural origin in 4-dimensional Chern-Simons theory that is compatible with the sigma-model variational principle, as will be reported elsewhere [11]. The key point is that ${\cal L}(\mathsf{x})$ is defined only up to gauge transformations, so the notion of a “reflected copy” should be formulated modulo this freedom. Accordingly we allow the reflected connection to be chosen within the gauge orbit of ${\cal L}(\rho(\mathsf{x}))$, and fix both the relevant representative $U$ and $\rho(\mathsf{x})$ by demanding that the divisor of zeros and poles on $\mathbb{CP}^{1}$ is preserved under the reflection. In the double-row picture the boundary values of this bulk gauge may be packaged as reflection matrices $K_{\sigma_{\star}}$, but the conceptual shift is that we use $U$ to determine $\rho(\mathsf{x})$ from bulk analyticity prior to solving for reflection data.
@@
As an important case, we apply this approach for string sigma-models on $AdS_{3}\times S^{3}\times M_{4}$ with mixed NSNS and RR fluxes; for reviews see [12, *Seibold:2024qkh]. The classical Lax connection of this model is explicitly known for all flux values [14] and illustrates the above limitation: the parity-based criterion alone does not identify $\rho(\mathsf{x})$ at nontrivial NSNS flux. Moreover, while worldsheet methods have led to $S$-matrices [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda], a dual CFT and lattice formulation remains absent for generic flux ratios. This makes the mixed-flux background an ideal arena to demonstrate our practical approach to boundary integrability.
@@
Beyond such technical advantages, boundary integrability of the mixed-flux $AdS_{3}\times S^{3}$ background is also motivated by exact worldsheet studies. At the pure NSNS point, the bosonic theory is described by an exactly solvable WZW model on $G=SU(1,1)\times SU(2)$ [20, *Maldacena:2000hw], in which CFT-preserving D-branes wrap twisted conjugacy classes that admit a boundary CFT (BCFT) description [22, *Stanciu:1999id, *Figueroa-OFarrill:2000gfl, 25]. Turning on RR-flux corresponds to a marginal deformation and, for pointlike branes, its effect on the open-string spectrum has been analysed by boundary conformal perturbation theory, while in the closed-string sector the RR perturbation is singular [26]. The boundary sector thus provides one of the few controlled windows on mixed-flux physics. It was in fact suggested in [14] that identifying integrable boundary conditions could provide the missing input at finite flux from integrability and offer a bridge to the BCFT picture. The present Letter realises this at the classical bosonic level. Imposing our analytic criterion on the mixed-flux Lax yields two branches of integrable boundary conditions: one compatible with the sigma-model action only in the pure RR limit, while the other persists for generic flux and includes all D-branes wrapping twisted conjugacy classes. Interestingly, at the level of the reflected copy, the former corresponds to identifying $\rho(\mathsf{x})$ from the invariance of ${\cal L}(\mathsf{x})$ under parity alone, whereas the latter requires combining worldsheet parity with a transformation within the Lax gauge orbit. It is in the latter branch that we have a rigid class of integrable branes whose embedding is fixed by the undeformed metric. We show that their gauge-invariant worldvolume two-form can remain flux-independent, while the mixed-flux contribution is encoded entirely in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ that are dynamical. [^3] This is consistent with the $\lambda$-deformed WZW case, where integrable D-branes also remain twisted conjugacy classes but for which the gauge-invariant worldvolume two-form acquires a non-trivial $\lambda$–dependence given the $\lambda$–deformed metric [27]. This provides a natural setup for comparing integrability-based reflection data with conformal perturbation theory around the solvable WZW point, where the same branes underlie the BCFT description. In particular, it would be very interesting to include the fermionic sector, study the open spectrum and connect to [26], which we leave for future work.
@@ I Analytic criteria
To analyse integrable two-dimensional sigma-models with boundaries, we start from bulk data: a flat Lax connection ${\cal L}(\mathsf{x})$, generally meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$ and in an $n\times n$ matrix representation of the loop algebra ${\mathbb{C}(\mathsf{x})}\otimes\mathfrak{g}$, whose analytic structure in $\mathsf{x}$ encodes an infinite tower of conserved charges. Under periodic conditions on $S^{1}_{\sigma}$, these are generated by $T_{{\cal L}}(2\pi,0;\mathsf{x})$. One can also allow for flatness-preserving gauge transformations
@@
with any dynamical $n\times n$ matrix $U:=U(\tau,\sigma)$. Then $T_{{\cal L}^{U}}(2\pi,0;\mathsf{x})=U(\tau,2\pi)T_{\cal L}(2\pi,0;\mathsf{x})U(\tau,0)^{-1}$ also generates conserved charges when $U$ is periodic in $\sigma$.
@@
which generates an infinite tower of conserved quantities under the integrable boundary conditions [^4] Flatness of the Lax and the boundary conditions (4) indeed imply that $\partial_{\tau}\mathrm{Tr}T_{\text{\tiny b}}(\mathsf{x})^{p}=0$ for any integer $p$, so that expanding the eigenvalues of $T_{\text{\tiny b}}(\mathsf{x})$ in $\mathsf{x}$ produces conserved objects.
@@
with $\overset{\text{\tiny b}}{=}$” denoting evaluation at the boundary $\sigma_{\star}=0,\pi$ and $K_{\sigma_{\star}}(\mathsf{x}):=K(\tau,\sigma_{\star};\mathsf{x})$ is the reflection matrix.
Typically, $K_{\sigma_{\star}}(\mathsf{x})$ is an $n\times n$ matrix in $\mathbb{C}(\mathsf{x})\otimes G$, with $G$ the Lie group of $\mathfrak{g}$. One may also allow constant matrices $K=w\not\in G$ that implement (possibly outer) automorphisms of $\mathfrak{g}$ as $W(y)=wyw^{-1}\in\mathfrak{g}$ for $y\in\mathfrak{g}$ [7]. At the lattice level, such choices can correspond to different representation assignments in the reflected monodromy. In suitable matrix realisations, they can also cover the common alternative convention where the inverse in (3) is replaced by transposition.
Note that in (4) only the endpoint values $U(\tau,\sigma_{\star})$ enter, and with (2) one may thus see the bulk gauge $U$ as part of a reflection matrix $\tilde{K}=KU$. However, as discussed earlier, the role of $U$ is fundamentally distinct: it is used to select a $\rho(\mathsf{x})$ for the bulk, while we reserve the notation $K$ for genuine, $\sigma$-independent, reflection matrices. This viewpoint will let us systemise the prescription of [4, 6, 7, 8, 28, 27, 29] through model-independent arguments that are not based on parity-invariant gluing.
Fixing $U$ and $\rho(\mathsf{x})$ from analyticity.
Our main point is that $\rho(\mathsf{x})$ can be fixed intrinsically from comparing the analytic structure of ${\cal L}(\mathsf{x})$ and the reflected representative ${\cal L}^{U}(\rho(\mathsf{x}))$, mirroring the lattice logic where $\rho(\mathsf{x})$ is derived from bulk properties of the $R$-matrix (rather than boundary $K$-matrices).
A first crucial ingredient is a choice of “physical” representatives for the unreflected and reflected rows, such that $T_{\text{\tiny b}}(\mathsf{x})$ can generate a set of local charges. Concretely, we require that there is a nonsingular $\mathsf{x}_{0}\in\mathbb{CP}^{1}$ where
${\cal L}_{\sigma}(\mathsf{x}_{0})=0,\qquad{\cal L}_{\sigma}^{U}(\rho(\mathsf{x}_{0}))=0,$ (5)
so that the double-row monodromy admits a local expansion around $\mathsf{x}_{0}$. Note that $K_{\sigma_{\star}}(\mathsf{x})$ is $\sigma$-independent and thus, when regular at $\mathsf{x}_{0}$, does not alter locality.
Let us now denote by $\mathrm{D}_{\text{\tiny p}}[{\cal L}(\mathsf{x})]$ and $\mathrm{D}_{\text{\tiny z}}[{\cal L}(\mathsf{x})]$ the pole and zero divisors of ${\cal L}(\mathsf{x})$ on $\mathbb{CP}^{1}$. Given (4) and (5), we will determine $\rho(\mathsf{x})$ from the requirement
@@
[…]
To determine $U$, one then scans for those gauges under which the zero divisor changes. Typically useful ones are those where the sigma-model fields are collected in a group element, $U=g(\tau,\sigma)\in G$. One can also include constant matrices $U=w\not\in G$ (cf. earlier), but these do not affect the divisor structure.
Furthermore, we consider $\rho:\mathbb{CP}^{1}\rightarrow\mathbb{CP}^{1}$ to be a reflection map that accompanies the spatial reflection, and hence it is an involutive Möbius transformation:
@@
with $\rho^{2}(\mathsf{x})=\mathsf{x}$ and $a,b,c\in\mathbb{C}$ to be determined. It can therefore only permute the marked points of (6) in pairs (or fix them), while preserving their analytic type. For a meromorphic Lax family with finitely many marked points under different gauges, this then leaves only a finite set of involutive $\rho(\mathsf{x})$ and $U$ candidates, that together with different $K_{\sigma_{\star}}$-solutions of (4) can lead to distinct integrable boundary conditions.

### v2

@@ Abstract
Boundary integrability provides rare analytic control over field theories with interfaces, from quantum impurity problems to open string dynamics. We propose an analytic approach for integrable boundaries in two-dimensional sigma-models that determines admissible reflection maps directly from the divisor structure of the Lax connection. Applied to open strings on $AdS_{3}\times S^{3}$ with mixed NSNS-RR flux, we find two branches of integrable boundaries: one restricted to pure RR flux, and another admitting D-branes wrapping twisted conjugacy classes for generic flux. At the WZW point, these reduce to the known conformal D-branes, opening a path to comparison with conformal perturbation theory. More broadly, our framework suggests generalisations of standard lattice constructions that may enlarge existing classifications of integrable boundaries.
@@
Field theories with interfaces play an important role in both high-energy and condensed-matter physics. In open string theory, worldsheet boundary conditions describe D-branes, while in condensed-matter and lattice systems they model boundary critical phenomena and impurities such as the Kondo effect. When the bulk theory is integrable, the situation becomes especially interesting, as an infinite set of conserved charges provides a rare source of analytic control for interacting dynamics. The presence of an interface probes this structure in a nontrivial way. Geometric symmetries are typically broken, and it is far from obvious that an infinite tower of conserved charges can survive. When they do, notions of boundary integrability have e.g. opened access to exact reflection data [1] and methods for computing defect observables in AdS/CFT [2]. Identifying which physical interfaces are compatible with integrable dynamics is therefore a basic structural question. In this Letter, we address it for two-dimensional sigma-models using simple analytical arguments from a novel perspective.
The starting point is an integrable sigma-model on $\mathbb{R}_{\tau}\times S^{1}_{\sigma}$ with an on-shell flat Lax connection ${\cal L}(\mathsf{x})$, [^1] For brevity, and where possible, we will suppress explicit time ($\tau\in\mathbb{R}_{\tau}:=\mathbb{R}$) or space ($\sigma\in S^{1}_{\sigma}:=S^{1}$ with radius 1) dependence. i.e. $\mathrm{d}{\cal L}(\mathsf{x})+{\cal L}(\mathsf{x})\wedge{\cal L}(\mathsf{x})=0$, that in general is meromorphic in $\mathsf{x}\in\mathbb{CP}^{1}$ and represents the equations of motion for every value of the spectral parameter $\mathsf{x}$. In this case, and under periodic boundary conditions ${\cal L}(\sigma+2\pi;\mathsf{x})={\cal L}(\sigma;\mathsf{x})$, the monodromy matrix $T_{\cal L}(2\pi,0;\mathsf{x})$, with
@@
On an interval $\sigma\in[0,\pi]$, boundary integrability requires a modification of this generator. The prototype is obtained by following Sklyanin’s construction [4, 5], where one forms a double-row monodromy matrix $T_{\text{\tiny b}}(\mathsf{x})$ by gluing $T_{\cal L}(\pi,0;\mathsf{x})$ over the interval $\sigma\in[0,\pi]$ to a reflected copy. On the lattice, Sklyanin’s approach applies to integrable models with crossing-symmetric $R$-matrices of difference form [5]. The reflection compatible with integrability then has two components: a spatial flip reversing the order of sites, and a spectral reflection $\mathsf{x}\mapsto\rho(\mathsf{x})=-\mathsf{x}$ implementing crossing symmetry. Additional boundary degrees of freedom are encoded in reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ at the endpoints $\sigma_{\star}=0,\pi$, which must satisfy the Boundary Yang-Baxter Equation (BYBE) to preserve integrability. The analogue in the sigma-model is to apply a worldsheet parity flip $\sigma\rightarrow 2\pi-\sigma$ combined with a spectral reflection $\rho(\mathsf{x})$ that ensures invariance of the Lax connection, see e.g. [4, 6, 7, 8]. One can then effectively set $T_{\text{\tiny b}}(\mathsf{x})=K_{0}(\mathsf{x})T_{\cal L}^{-1}(\pi,0;\rho(\mathsf{x}))K_{\pi}^{-1}(\mathsf{x})T_{\cal L}(\pi,0;\mathsf{x})$ and solve the conditions for $K_{\sigma_{\star}}$ under which infinite charges are generated by $T_{\text{\tiny b}}(\mathsf{x})$.
@@
However, there are limitations to taking this prescription directly as a definition of boundary integrability for sigma-models. On the lattice, $\rho(\mathsf{x})$ is dictated by the crossing structure of a chosen bulk $R$-matrix. For an integrable sigma-model, such input is often unavailable (or indirect through the sigma-model $S$-matrix), and the defining object is just a Lax connection ${\cal L}(\mathsf{x})$. In addition, there need not exist a $\rho(\mathsf{x})$ that leaves ${\cal L}(\mathsf{x})$ invariant after a parity flip. In these cases, there is no given principle to select an appropriate generator of conserved charges.
@@
In this Letter, we will instead bypass direct input on bulk parity or crossing invariance and determine $\rho(\mathsf{x})$ from the analytic structure of the Lax connection. [^2] For alternative approaches via parity, yet with additional assumptions, see [9, *Corrigan:1996nt]. Here we will take an analytic approach that moreover admits a natural origin in 4-dimensional Chern-Simons theory, as will be reported elsewhere [11]. A central shift in our approach is to fix $\rho(\mathsf{x})$ by treating the reflected copy of ${\cal L}(\mathsf{x})$ modulo gauge transformations that preserve the flatness condition. Our prescription to fix both $\rho(\mathsf{x})$ and the gauge representative $U$ is then to simply demand that the divisor of zeros and poles of ${\cal L}(\mathsf{x})$ on $\mathbb{CP}^{1}$ is preserved under the reflection. Once done, residual boundary reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$ can be introduced as above to select distinct integrable interfaces.
@@
As an important case, we apply this approach to the string sigma-model on $AdS_{3}\times S^{3}\times M_{4}$ with mixed NSNS and RR flux; for reviews see [12, *Seibold:2024qkh]. Its classical Lax connection is known for all flux values [14] and illustrates the above limitation: the parity-based criterion alone does not fix $\rho(\mathsf{x})$ at nontrivial NSNS flux. Moreover, while worldsheet methods have led to $S$-matrices [15, *Lloyd:2014bsa, *OhlssonSax:2023qrk, *Frolov:2025uwz, *Frolov:2025tda], a dual CFT and lattice formulation remains absent for generic fluxes. This makes the mixed-flux model an ideal setting to demonstrate our approach to boundary integrability.
@@
Boundary integrability of the mixed-flux $AdS_{3}\times S^{3}$ background is also inherently motivated by exact worldsheet studies. At the pure NSNS point, the bosonic theory is described by an exactly solvable WZW model on $SU(1,1)\times SU(2)$ [20, *Maldacena:2000hw], in which CFT-preserving D-branes wrap twisted conjugacy classes that admit a boundary CFT (BCFT) description [22, *Stanciu:1999id, *Figueroa-OFarrill:2000gfl, 25]. Turning on RR-flux corresponds to a marginal deformation and, for pointlike branes, its effect on the open-string spectrum has been analysed through boundary conformal perturbation theory, leading to all-orders resummed results, while in the closed-string sector the RR perturbation is singular [26]. The boundary sector thus provides one of the few controlled windows on mixed-flux physics. It was in fact suggested in [14] that identifying integrable boundary conditions could provide further structural information at finite flux and offer a bridge to the BCFT picture. The present Letter realises this at the classical bosonic level. Imposing our analytic criterion on the mixed-flux Lax, we find two branches of integrable boundary conditions: one compatible with the sigma-model action only in the pure RR case, while the other persists for generic flux and includes all D-branes wrapping twisted conjugacy classes. Interestingly, the former arises by identifying $\rho(\mathsf{x})$ from the invariance of ${\cal L}(\mathsf{x})$ under parity alone, whereas the latter requires combining parity with a gauge transformation of the Lax connection. It is in the second branch that we have a rigid class of integrable branes whose embedding coincides with the conformal WZW branes. We show that their worldvolume two-form can remain flux-independent, with the mixed-flux contribution encoded entirely in dynamical reflection matrices $K_{\sigma_{\star}}(\mathsf{x})$. [^3] This is compatible with the $\lambda$-deformed WZW case, where integrable D-branes also remain twisted conjugacy classes but for which the gauge-invariant worldvolume two-form acquires a non-trivial $\lambda$–dependence given the $\lambda$–deformed metric [27].
Outlook. For mixed-flux $AdS_{3}\times S^{3}$, our results thus provide a natural setup to compare integrability-based data with conformal perturbation theory around the solvable WZW point, where the same twisted conjugacy classes underlie the BCFT description. Including the fermionic sector and analysing the open string spectrum would enable comparison with the all-order spectra of [26] and open the way to extending the analysis beyond the pointlike brane sector. We leave this for future work.
More broadly, our approach not only generalises sigma-model constructions but also suggests a possible generalisation for lattice models. Traditionally, the BYBE assumes that both rows in the double-row monodromy share the same $R$-matrix. By allowing different representatives in each row, mirroring different gauge choices of ${\cal L}(\mathsf{x})$, a natural question is whether classifications can be enlarged to include new integrable interfaces that yield distinct states after reflection. Given the scarcity of integrable models with interfaces—for defect observables in holography, benchmarking impurity physics, engineering quantum circuits, etc.—this could provide valuable new ground for applications of exact methods.
@@ I Analytic criteria
To analyse integrable two-dimensional sigma-models with boundaries as outlined in the introduction, we start from bulk data: an on-shell flat Lax connection ${\cal L}(\mathsf{x})$, generally in an $n\times n$ matrix representation of the loop algebra ${\mathbb{C}(\mathsf{x})}\otimes\mathfrak{g}$. Under periodic conditions on $S^{1}_{\sigma}$, towers of charges are then generated by $T_{{\cal L}}(2\pi,0;\mathsf{x})$. One can also allow for flatness-preserving gauge transformations
@@
with any dynamical $n\times n$ matrix $U:=U(\tau,\sigma;\mathsf{x})$. Then $T_{{\cal L}^{U}}(2\pi,0;\mathsf{x})=U(\tau,2\pi;\mathsf{x})T_{\cal L}(2\pi,0;\mathsf{x})U(\tau,0;\mathsf{x})^{-1}$ is also a generator of charges when $U$ is periodic in $\sigma$.
@@
which generates an infinite tower of conserved charges under the integrable boundary conditions [^4] Flatness of the Lax and the boundary conditions (4) indeed imply that $\partial_{\tau}\mathrm{Tr}T_{\text{\tiny b}}(\mathsf{x})^{p}=0$ for any integer $p$, so that expanding the eigenvalues of $T_{\text{\tiny b}}(\mathsf{x})$ in $\mathsf{x}$ produces conserved objects.
@@
with $\overset{\text{\tiny b}}{=}$” denoting evaluation at the boundary $\sigma_{\star}=0,\pi$ and $K_{\sigma_{\star}}:=K(\tau;\mathsf{x})$ $\sigma$-independent reflection matrices.
Remarks: (i) Constant matrices $K_{\sigma_{\star}}$ or $U=w$, can implement automorphisms of $\mathfrak{g}$ when $W(y)=wyw^{-1}\in\mathfrak{g}$ for $y\in\mathfrak{g}$ [7]. At the lattice level, these can give different representation choices for the reflected bulk sites. In suitable realisations, this covers the common alternative convention where the inverse in (3) is replaced by transposition. (ii) Only the endpoint values $U(\tau,\sigma_{\star})$ enter in (4), and given (2) one may thus see the bulk gauge $U$ as part of a boundary matrix $\tilde{K}=KU$. However, for us the role of $U$ is fundamentally distinct: it will determine relevant $\rho(\mathsf{x})$ for the reflected bulk region, prior to solving for boundary data. In other words, $K_{\sigma_{\star}}$ is the residual gauge freedom at the boundary after reflection. This viewpoint will let us generalise the prescription of [4, 6, 7, 8, 28, 27, 29] beyond parity-invariant gluing.
Fixing $U$ and $\rho(\mathsf{x})$ from bulk analyticity.
Our main point is that $U$ and $\rho(\mathsf{x})$ can be fixed intrinsically from the analytic structure of ${\cal L}(\mathsf{x})$, mirroring the lattice logic where $\rho(\mathsf{x})$ is derived from bulk properties of the $R$-matrix (rather than boundary $K$-matrices).
A first useful criterion is a choice of “physical” representatives for the unreflected and reflected rows, such that $T_{\text{\tiny b}}(\mathsf{x})$ can generate a set of local charges. Concretely, one requires that there exist points $\mathsf{x}_{0},\mathsf{x}^{\prime}_{0}\in\mathbb{CP}^{1}$ where
${\cal L}_{\sigma}(\mathsf{x}_{0})=0,\qquad{\cal L}_{\sigma}^{U}(\mathsf{x}^{\prime}_{0})=0,$ (5)
so that $T_{\text{\tiny b}}(\mathsf{x})$ has a local expansion around $\mathsf{x}_{0}$ when $\mathsf{x}^{\prime}_{0}=\rho(\mathsf{x}_{0})$ for some $\rho(\mathsf{x})$. [^5] Note that $K_{\sigma_{\star}}$ is $\sigma$-independent and thus, when regular at $\mathsf{x}_{0}$, does not alter locality. In practice, relevant representatives are found by scanning for those gauges under which ${\cal L}^{U}_{\sigma}$ remains local while its zero divisor changes. Note, however, that this is a strong requirement: a local $U(\mathsf{x}_{0}^{\prime})$ where ${\cal L}_{\sigma}^{U}(\mathsf{x}^{\prime}_{0})=0$ requires ${\cal L}_{\sigma}(\mathsf{x}_{0}^{\prime})$ to be off-shell flat. Relaxing these criteria would formally enlarge the possibilities, but could either introduce incompatibilities with the conditions (4) for trivial $K_{\sigma_{\star}}$ or obstruct their interpretation as local boundary actions as e.g. in (22).
[…]
@@
Furthermore, we consider $\rho:\mathbb{CP}^{1}\rightarrow\mathbb{CP}^{1}$ to be a nontrivial reflection map that accompanies the spatial reflection, and hence as the involutive Möbius transformation:
@@
with $\rho^{2}(\mathsf{x})=\mathsf{x}$ and $a,b,c\in\mathbb{C}$ to be determined. It can therefore only permute the marked points of (6) in pairs (or fix them), while preserving their analytic type. For a meromorphic Lax connection with finitely many marked points under different gauges, this then leaves only a finite set of possible $\rho(\mathsf{x})$ and $U$ candidates, that together with different $K_{\sigma_{\star}}$ solutions of (4) can lead to distinct integrable boundary conditions.